if sin4A/a + cos4B/b = 1/(a+b) then prove that sin8A/a3 + cos8B/b3 =1/(a+b)3

Grade:12th pass

1 Answers

Lab Bhattacharjee
8 years ago
\text{HINT: I believe the question to be} \\ \dfrac{\sin^4A}a+\dfrac{\cos^4A}b=\dfrac1{a+b} \\ \text{Set }\sin^2A=x\implies\cos^2x=1-x\text{ to find } x=\dfrac a{a+b}